Circular No. 7024

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1998ec IN UGC 3576 S. Jha, P. Garnavich, P. Challis, and R. Kirshner, Harvard-Smithson-

ian Center for Astrophysics, report that a spectrum of SN 1998ec, taken by P. Berlind on Sept. 29.5 UT with the F. L. Whipple Observatory 1.5-m telescope (+ FAST spectrograph), reveals it to be of type Ia before maximum light. The NASA/IPAC Extragalactic Database recession velocity for UGC 3576 is 5948 km/s, which implies an expansion velocity of 13 000 km/s for the supernova atmosphere, derived from the strong Si II 635.5-nm absorption feature.

NGC 7582 B. Joguet, European Southern Observatory (ESO); D. Kunth, Insti-

tut d’Astrophysique de Paris; and R. Terlevich, Institute of Astronomy, Cambridge, report the discovery of a bright flare in the Seyfert 2 galaxy, NGC 7582, as revealed by spectroscopy (integration time 2700 s; spectral resolution 0.3 nm; range 370–670 nm) performed on July 11 with the ESO 1.54-m telescope at La Silla: “The galaxy is located at α = 23^{h}18^{m}23^{s}.50, δ = −42^{o}22′14′′.0 (equinox 2000.0). From the spectrum, broad line components at Hα and Hβ are seen with a blueshift of ∼ 2500 km/s with respect to the narrow lines. The FWZI of the broad Hα component is ∼ 25 000 km/s. Also, broad Fe II emission lines can be observed around 440 and 540 nm. Another spectrum (integration time 3600 s; spectral resolution 0.2 nm; range 340–540 nm) taken on June 1 with the ESO 1.52-m telescope does not show any sign of broad lines in the nucleus of this galaxy. The continuum brightened by about a factor of two at 540 nm between the June and July spectra. The nuclear magnitude computed with the July spectrum at 540 nm is 14.3.”

COMET P/1998 QP_{54} (LONEOS-TUCKER)

Improved orbital elements from MPC 32595:

T = 1998 Oct. 6.4196 TT  ω = 30.^{o}1103
e = 0.551814  Ω = 341.9272 }2000.0
q = 1.881718 AU  i = 17.7418

a = 4.198518 AU n^{o} = 0.1145671 P = 8.603 years

1998 October 6  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green